Test for exactness and find the complete solution of y/(x-1) dx+ {ln(2x-2)+ 1/y}dy=0

1
Expert's answer
2022-03-03T08:51:22-0500

We check the differential equation for exactness:


"\dfrac{y}{x-1}dx+(\ln(2x-2)+\dfrac{1}{y})dy=0"

"M(x,y)=\dfrac{y}{x-1}, \dfrac{\partial M}{\partial y}=\dfrac{1}{x-1}"

"N(x, y)=\ln(2x-2)+\dfrac{1}{y}, \dfrac{\partial N}{\partial x}=\dfrac{1}{x-1}"

"\dfrac{\partial M}{\partial y}=\dfrac{1}{x-1}= \dfrac{\partial N}{\partial x}"

We see that "\dfrac{\partial M}{\partial y}= \dfrac{\partial N}{\partial x}," so that this equation is exact.


We have the following system of differential equations to find the function "u(x, y)"


"\dfrac{\partial u}{\partial x}=\dfrac{y}{x-1}"

"\dfrac{\partial u}{\partial y}=\ln(2x-2)+\dfrac{1}{y}"

Hence


"u(x, y)=\int\dfrac{y}{x-1}dx+\varphi(y)"

"=y\ln(x-1)+\varphi(y)"

Now, by differentiating this expression with respect to "y" and equating it to "\dfrac{\partial u}{\partial y}," we find the derivative "\varphi'(y)"


"\ln(x-1)+\varphi'(y)=\ln(2x-2)+\dfrac{1}{y}"

"\varphi'(y)=\ln2+\dfrac{1}{y}"

"\varphi(y)=\int (\ln2+\dfrac{1}{y})dy=y\ln2+\ln y-C"

Thus, the general solution of the differential equation is


"y\ln(2x-2)+\ln y=C"